OpenStudy (anonymous):

What quantity of energy (in kJ) is contained in a burst of gamma rays which contains 1.75 moles of photons which have a frequency of 5.80 * 1014 Hz? Can some help me solve this step by step?

OpenStudy (anonymous):

The energy of a photon is given by this equation: \[E = h * \nu\] Where nu is the frequency and h is Planck's constant: \[h = 6.626 *10^{-34} J*s\] Therefore, the energy of each of our gamma ray photons is: \[E = (6.626 * 10^{-34} J*s)(5.80*10^{14} s ^{-1}) = 38.4308*10^{-20} J\] As we also know, there are 6.022e23 of anything in a mole, so we have 10.5385e23 photons. Multiplying this by the energy for one photon gives us the total energy: \[38.4308*10^{-20} J * 10.5385*10^{23} = 405.00298*10^{3} J = 405.00298 kJ\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Sure! Glad to help.